OpenStudy (anonymous):

Marelle bought a total of 6 drinks and 6 bags of popcorn for herself and five friends at the ballpark. She spent a total of $23.70. Part 1: Write the equation for this word problem. Be sure to identify your variables. Part 2: If each bag of popcorn cost $2.25, solve the equation to find the cost of each drink. Show your work.

OpenStudy (anonymous):

1. Let's say variable p indicates the price of one bag of popcorn and variable d indicates the price of one drink. She bought 6 of each, so you would have 6p and 6d. Finally, this totaled 23.70, so the equation would be 6p+6d=23.70 2. Now, popcorn costs 2.25. Replace the variable for the cost of popcorn in the equation with 2.25, giving you 6(2.25)+6d=23.70. Then, solve: 6(2.25)+6d=23.70 (Multiply 6 by 2.25) 13.5+6d=23.70 (Substract 13.5 from both sides) 6d=10.2 (Divide both sides by 6) d=1.7 Each drink costs $1.70